OpenStudy (anonymous):

Solve 3(x + 1) + 6 = 33

pooja195 (pooja195):

Use the distributive property 3(x + 1) What do you get once you distribute the 3?

OpenStudy (jacob902):

Distribute the 3. 3x + 3 + 6 = 33 Isolate the x term. 3x = 33 - 3 - 6 Simplify. 3x = 24 Divide. x = 8 Check answer. 3[(8) + 1] + 6 = 33 3[9] + 6 = 33 27 + 6 = 33 33 = 33

pooja195 (pooja195):

Or just give the answer .-.

OpenStudy (anonymous):

3(x)+3(1) lol

pooja195 (pooja195):

good 3x+3+6=33 Combine like terms

pooja195 (pooja195):

6+3= ?

OpenStudy (anonymous):

9

pooja195 (pooja195):

3x+9=33 subtract 9 from both sides because we are triny to isolate x 9-9=0 33-9=?

pooja195 (pooja195):

*trying

OpenStudy (anonymous):

lol 24

pooja195 (pooja195):

3x=24 divide by 3 from both sides 3/3=1 so its only x x=24/3= ?

OpenStudy (anonymous):

8

pooja195 (pooja195):

Good so we get x=8 :)

OpenStudy (anonymous):

so x=8 is my answer right? :)

pooja195 (pooja195):

yes! :)

OpenStudy (anonymous):

thnx :)

pooja195 (pooja195):

No problem :D